OpenStudy (anonymous):

Which two points are on the graph of T(b) = –20 – 5b? Choose exactly two answers that are correct. (Points : 1) (–4, 0) (4, 0) (0, –20) (0, 20)

OpenStudy (anonymous):

@Karon86

OpenStudy (karon86):

hi

OpenStudy (anonymous):

hello. do you know how to do it?

OpenStudy (karon86):

I will try

OpenStudy (karon86):

I am catching on

OpenStudy (karon86):

What type of function is this

OpenStudy (karon86):

Is it f(x)

OpenStudy (anonymous):

yea

OpenStudy (karon86):

wow you are in luck because i knows this very well

OpenStudy (anonymous):

yahh!

OpenStudy (karon86):

okay so f(x)=y

OpenStudy (karon86):

That is one formula

OpenStudy (karon86):

The reason why f(x)=y because it is written as a function rule

OpenStudy (anonymous):

what?

OpenStudy (karon86):

T(b) is really f(x) its just writen as T(b) . This whole equation is a function. f(x) is y y is part of y point on (x,y) x is a part of the x point on (x,y)

OpenStudy (karon86):

So T(b)=-20-5b is really f(x)=-20-5x

OpenStudy (karon86):

Do you get it?

OpenStudy (karon86):

But the proble is not suppose to be in that form in order to find the points because the type of equation is a linear equation

OpenStudy (karon86):

A linear equation is written in y=mx+b and b is the y-intercept, m is the slope. So you have to find the slope and the y-int

OpenStudy (anonymous):

so what r the anseres?

OpenStudy (karon86):

T(b)=-20-5b has to be put inthe form of y=mxtb So lets rearrange it: T(b)=-20-5b T(b)=-5b-20 by the commutative property (a+b=b+a) this is your equation but you have to put it in the calculator as y=-5-20

OpenStudy (karon86):

You should find your points on the calculator by entering the linear equation

OpenStudy (karon86):

I will give you one of the points which is (0,-20)

OpenStudy (karon86):

Type if you are with me

OpenStudy (anonymous):

i am now sorry my friend fell and got hurt

OpenStudy (anonymous):

is the other one 0,-4

OpenStudy (anonymous):

tomorro will you be able to help me with a few more?? @Karon86

OpenStudy (karon86):

hold on a sec

OpenStudy (karon86):

no

OpenStudy (karon86):

Did you graph it?

OpenStudy (karon86):

Go on graphing calculator online. Remember, (x,y) is your formula for a point

OpenStudy (karon86):

Oh i am so sorry. I will give you the answer (-4,0)

OpenStudy (anonymous):

thank you :) i wil.l tag you in my qyestions tomoro

OpenStudy (anonymous):

I will tag you in my questions tomorrow*****